# Kernel Banzhaf: A Fast and Robust Estimator for Banzhaf Values

### v1

@@ Appendix G Kernel Banzhaf with Sampling without Replacement
Figure 10: Comparison of $\ell_{2}$-norm error between Kernel Banzhaf and Kernel Banzhaf with sampling without replacement (SWOR) across eight datasets. When sample size $m<2^{n}$, the performance of the two algorithms is quite similar, as illustrated by the overlapping lines. Note that the sharp error reduction in the first three small datasets occurs when sample size $m$ exceeds $2^{n}$, due to the SWOR algorithm exhaustively sampling all possible subsets at $m>2^{n}$. Theoretically, our sampling with replacement approach is better in terms of efficiency and its use of exact $m$ samples.
Contrary to the approach in Musco & Witter (2024), which employs sampling without replacement to enhance KernelSHAP, our empirical results, as illustrated in Figure 10, indicate that this method does not yield improvements. We observe a significant reduction in error when the sample size $m$ exceeds $2^{n}$, where $n$ is the number of features, in the first three small datasets. This reduction can be attributed to the fact that the algorithm has exhaustively sampled all possible subsets once $m>2^{n}$. However, it is impractical to set $m$ beyond the total number of subset combinations $2^{n}$ in typical scenarios, as doing so would allow a complete enumeration of subsets, thereby directly computing the ground truth Banzhaf values.
In standard cases where $m<2^{n}$, our sampling with replacement approach not only matches the performance of sampling without replacement but also surpasses it in terms of efficiency and scalability. This is evident as sampling using the binary representation of random integers used in sampling without replacement becomes infeasible for large $n$. Additionally, sampling with replacement ensures the collection of exactly $m$ distinct samples, in contrast to the approximate number $m^{\prime}$ obtained without replacement. Given these advantages, we continue to utilize the sampling without replacement setting in our proposed algorithm, offering a comprehensive analysis that diverges from the findings presented in Musco & Witter (2024).
Appendix H Properties of Shapley Values and Banzhaf Values
Shapley values satisfy four desirable properties: Null Player ensures that a player who does not contribute to any coalition, meaning their inclusion in any subset of players does not affect the overall outcome, is assigned a value of zero. Symmetry requires that two players who contribute equally to all possible coalitions receive the same value. Linearity requires that the Shapley value of a player in a combined game (formed by adding two games together) is equal to the sum of that player’s Shapley values in the two individual games. Efficiency requires that the total value assigned to all players must sum to the value generated by the full set of players. (Shapley, 1953).
Instead of the Efficiency property, the Banzhaf index satisfies 2-Efficiency, which requires that the sum of the values of any two players equals the value of these two players when considered jointly in a reduced game setting (Banzhaf, 1965; Lehrer, 1988). The necessity of the efficiency property has been debated in the context of machine learning. Sundararajan et al. (2017) suggest that the Efficiency property is only essential in contexts where semi-values, such as those in voting games, are interpreted numerically; Kwon & Zou (2022) argues that the utility function in machine learning applications often does not correspond directly to monetary value, so aligning the sum of data values with total utility is unnecessary. In applications where the primary goal involves ranking features according to their importance or evaluating data, the exact numerical contribution of each feature is less critical. Both Banzhaf and Shapley values, despite their theoretical disparities, often yield the same ordering of players as shown in Karczmarz et al. (2022), which suffices for these applications. Therefore, given their efficiency and robustness properties, Banzhaf values serve as particularly effective tools in machine learning tasks (Wang & Jia, 2023).

### v2

@@ Appendix G Ablation: Kernel Banzhaf With and Without Replacement
Figure 8: Comparison of relative squared $\ell_{2}$-norm error between Kernel Banzhaf and Kernel Banzhaf with sampling without replacement (SWOR) across eight datasets. When sample size $m<2^{n}$, the performance of the two algorithms is quite similar, as illustrated by the overlapping lines. Note that the sharp error reduction in the first three small datasets occurs when sample size $m$ exceeds $2^{n}$, due to the SWOR algorithm exhaustively sampling all possible subsets at $m>2^{n}$.
Appendix H Feature Ranking Recovery
Aside from evaluating the quantitative errors between estimated Banzhaf values and exact Banzhaf values, another critical metric that reveals the meaningfulness of the estimated results is how well the estimator recovers feature ranking. Feature ranking is important to feature comparison and selection, which are useful for enhancing the performance of machine learning models. Accurate feature ranking helps in identifying the most influential features, thereby facilitating more efficient and effective feature engineering and dimensionality reduction strategies.
In order to evaluate this property, we incorporate two well-known metrics: Cayley distance and Spearman rank correlation. The Cayley distance refers to the minimum number of transpositions required to transform one permutation into another. This metric provides a concrete measure of the difference between two rankings, capturing the minimal edit sequence needed, which is particularly useful in understanding the stability and reliability of feature ranking methods, and it’s also adopted in Karczmarz et al. (2022). Spearman’s rank correlation, $\rho$, on the other hand, measures the strength and direction of association between two ranked variables. Formally, it is defined as the Pearson correlation coefficient between the rank values of the variables, mathematically expressed as:
$\rho=1-\frac{6\sum d_{i}^{2}}{n(n^{2}-1)}$
where $d_{i}$ represents the difference between the ranks of corresponding variables $x_{i}$ and $y_{i}$, and $n$ is the number of observations. This metric offers insights into how well the ranking produced by the estimator preserves the monotonic relationship compared to the exact ranking, providing a measure of ranking fidelity.
In scenarios with a large feature space, the most significant features often have a more pronounced impact on model predictions. In these cases, the overall ranking may be cluttered with a large number of features that show only minor differences in their Banzhaf values, making it difficult to distinguish among lower-ranked features effectively. Focusing on the top 20% of features, therefore, targets those variables most likely to affect predictive accuracy and model stability, offering a more pragmatic evaluation of ranking recovery.
In our evaluation, we identify the top 20% of features ranked by exact Banzhaf values. For any feature missing from the top 20% as derived from the estimated Banzhaf values, we add it to the end of the permutation. This adjustment ensures that the evaluation penalizes discrepancies at the top of the distribution, which are most critical for decision-making and model interpretation.
Our findings for the top 20% feature ranking recovery are presented in Figures 4 and 9. This task, focusing only on the ranks of the most significant features, is more practical. For datasets with feature size $\leq 20$, we take the top 7 features instead. Our Kernel Banzhaf algorithm consistently outperforms or matches the MC method across all datasets, demonstrating its effectiveness in identifying and ranking important features.
We also explore feature ranking recovery on all the features in Figures 10 and 11. MC gives the best performance, followed by Kernel Banzhaf. Because of its sampling structure i.e., $v(S\cup\{i\})-v(S)$, MC is uniquely suited to recover Banzhaf values near 0. As shown in Figure 12, there are many Banzhaf values near 0 and hence MC gives the best performance when recovering these features.
Figure 9: Comparison of top 20% feature ranking recovery using Spearman correlation (higher is more accurate). Kernel Banzhaf consistently gives the best performance.
Figure 10: Comparison on full feature ranking recovery using Cayley distance (lower is more accurate). MC gives the best performance, followed by Kernel Banzhaf. Because of its sampling structure i.e., $v(S\cup\{i\})-v(S)$, MC is uniquely suited to recover Banzhaf values near 0. As shown in Figure 12, there are many Banzhaf values near 0.
Figure 11: Comparison on full feature ranking recovery using Spearman correlation (higher is more accurate). MC gives the best performance, followed by Kernel Banzhaf. Because of its sampling structure i.e., $v(S\cup\{i\})-v(S)$, MC is uniquely suited to recover Banzhaf values near 0. As shown in Figure 12, there are many Banzhaf values near 0.
Figure 12: There are many Banzaf values near 0, giving MC an advantage if full feature ranking recovery.
Appendix I Adversarial Perturbations
To further demonstrate the robustness of our proposed estimator, instead of independently perturbing all queries to the set function, we only perturb sets $S$ that contain a chosen item $i$. In the first experiment setting, we select $i$ uniformly at random. Then, instead of observing $v(S)$ on the query to subset $S$, the algorithms observe $v(S)+x$ where $x$ 0 if $i\not\in S$ and, if $i\in S$, we have $x\sim\mathcal{N}(0,\sigma^{2})$ as before. These results appear in Figure 13.
Figure 13: Plots of relative squared $\ell_{2}$-norm error by adversarial noise levels across Banzhaf estimators. For each noise level $\sigma$, the estimator observes $v(S)+x$ where $x\sim N(0,\sigma)$ when set $S$ contains a randomly chosen feature $i$. Similar to previous experiments, Kernel Banzhaf outperforms for lower noise levels, eventually matching its ablated version and MSR for larger noise.
Beyond this, in a follow-up experiment, we introduce a more adversarial noise setting. Each algorithm is run once on set function $v$ (no perturbation in the query access). Then we compute the relative error of each estimated value $\tilde{\phi}_{j}$ relative to the baseline $\phi_{j}$. We select the item $i$ with the largest relative error. Then, we evaluate each algorithm as before, but now the queries are perturbed if the set $S$ contains the adversarially chosen $i$. These results appear in Figure 14.
Figure 14: Plots of relative squared $\ell_{2}$-norm error by adversarial noise levels across Banzhaf estimators. For each noise level $\sigma$, the estimator observes $v(S)+x$ where $x\sim N(0,\sigma)$ when set $S$ contains feature $i$ with the largest relative error in the normal setting. Similar to previous experiments, Kernel Banzhaf outperforms for lower noise levels, eventually matching its ablated version and MSR for larger noise.